OpenStudy (anonymous):

Given the point (5, 6) and the slope of 5, find y when x = 21.

OpenStudy (hoblos):

let's find the equation of the line first, then we can get y for any x can you find the equation ?

OpenStudy (anonymous):

y=5(21)+b ??

OpenStudy (hoblos):

yeah that's what you have to do to find y at x=21 but first you have to find the equation passing by (5,6) and of slope 5

OpenStudy (anonymous):

no idea lol

OpenStudy (hoblos):

apply same formula as previous question y - y1 = m( x - x1)

OpenStudy (anonymous):

y-6=5(x-5)

OpenStudy (hoblos):

correct now can you simplify it ?

OpenStudy (anonymous):

y-6=5x-25 y=5x-31